OpenStudy (anonymous):

Need help with integration. the integral of cos(ax)sin(bx)dx by using the cyclic trick. I cannot use the product to sum formula. @Mathematics

hero (hero):

You can use integration by parts

OpenStudy (amistre64):

\[\frac{cos((b-a))x}{2(b-a)}-\frac{cos((b+a)x)}{2(b+a)}\]

OpenStudy (amistre64):

i gots no idea what a cyclic trick is tho

hero (hero):

It's when you get the same integral on both sides when using integration by parts and then you add it to both sides. Something like that the OP was saying

OpenStudy (amistre64):

ohh ....

OpenStudy (amistre64):

we just did a set of stuff like this for a group project in class which is why i remember the formula ...

OpenStudy (amistre64):

or at least i hope i remembered it lol

OpenStudy (amistre64):

\[\int x^n\ ln(x)\ dx=\frac{x^{n+1}}{n+1}ln(x)-\frac{x^{n+1}}{2(n+1)}+C\] i think

OpenStudy (amistre64):

nah, thats (n+1)^2 , i knew there was a 2 in it

OpenStudy (anonymous):

The integration by parts trick goes as follows... \[ I = \int cos(ax)sin(bx) dx =\frac{- cos(ax)cos(bx)}{b} - \int\frac{a*sin(ax)cos(bx)}{b}dx \] \[ = \frac{-cos(ax)cos(bx)}{b} - \frac{a*sin(ax)sin(bx)}{b^2} + \int\frac{a^2cos(ax)sin(bx)}{b^2}dx\] \[= \frac{-cos(ax)cos(bx)}{b} - \frac{a*sin(ax)sin(bx)}{b^2} + \frac{a^2}{b^2}I \] so \[(1-\frac{a^2}{b^2})I = \frac{-cos(ax)cos(bx)}{b} - \frac{a*sin(ax)sin(bx)}{b^2} \] and, being careful to recognize that we need some constant C, \[I = \frac{ \frac{-cos(ax)cos(bx)}{b} - \frac{a*sin(ax)sin(bx)}{b^2}}{1-\frac{a^2}{b^2}} +C \] or equivalently, \[ I = \frac{b*cos(ax)cos(bx) + a*sin(ax)sin(bx)}{a^2-b^2} +C \]

hero (hero):

Someone clearly has a lot of free time on their hands

OpenStudy (anonymous):

It's either this or revising my statement of purpose for the billionth time...